# The ECME Algorithm Using Factor Analysis for DOA Estimation in Nonuniform Noise

### v1

@@
Abstract
Maximum likelihood factor analysis has been used for direction of arrival estimation in unknown nonuniform noise and some iterative approaches have been developed. In particular, the Factor Analysis for Anisotropic Noise (FAAN) method proposed by Stoica and Babu has excellent convergence properties. In this letter, the Expectation/Conditional Maximization Either (ECME) algorithm, an extension of the expectation-maximization algorithm, is designed, which has almost the same computational complexity at each iteration as the FAAN method. However, numerical results show that the ECME algorithm yields faster stable convergence and is computationally more efficient.
Index Terms:
Array signal processing, direction of arrival (DOA) estimation, expectation-maximization (EM) algorithm, factor analysis, maximum likelihood.
I Introduction
@@
Factor analysis (FA) has been a tool in psychology for over one hundred years [1] and its purpose is to extract or estimate the low-rank covariance matrix from a measurement covariance matrix. To this end, several techniques have been developed, but the only one available, which provides efficient and consistent estimators, is maximum likelihood FA (MLFA) [2]. Unfortunately, these estimators cannot be expressed in closed form, which leads to a variety of iterative approaches [3]. In particular, one of the most promising is the Expectation-Maximization (EM) algorithm [4].
@@
The EM algorithm regards observations or measurements as the incomplete data and designs the complete data containing missing (latent or unobserved) data. The complete data is so manipulable that its log-likelihood function (LLF) is more tractable than the incomplete-data or actual LLF. Hence, the EM algorithm uses the complete-data LLF to efficiently update the estimates of parameters, which can still increase the actual likelihood. To be specific, each iteration of the EM algorithm consists of an expectation step (E-step) and a maximization step (M-step). However, the M-step requires maximizing the complete-data LLF, which causes two shortcomings: 1) slow convergence, and 2) difficult maximization in numerous complex cases. To make up for both shortcomings, the Expectation/Conditional Maximization Either (ECME) algorithm, an extension of the EM algorithm, has been proposed [5] and then applied to MLFA [6]. However, a conditional M-step (CM-step) at each iteration utilizes Newton-Raphson to update the estimates of partial parameters by the actual LLF and thus is complex as well as unstable.
@@
In addition to psychometrics, econometrics, statistics and so on, FA recently finds applications in array signal processing, especially including direction of arrival (DOA) estimation in unknown nonuniform noise [7, 8]. The nonuniform noise assumption can let the signal received at an array be modeled by FA so that MLFA can be utilized to estimate the signal subspace for the subspace based DOA estimation [9]. For this purpose, an iterative approach, called Iterative ML Subspace Estimation (IMLSE), is proposed in [7]. This method sequentially updates the estimates of parameters at each iteration by explicit formulas and is computationally attractive. However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [8]. Then, they propose a new iterative approach, called Factor Analysis for Anisotropic Noise (FAAN), with excellent convergence properties.
@@
In this letter, the ECME algorithm is designed again for MLFA, which has almost the same computational complexity at each iteration as the FAAN method. However, numerical results show that the ECME algorithm yields faster stable convergence and is computationally more efficient. Importantly, the signal subspace estimated by the ECME algorithm can be used for the subspace based DOA estimation in unknown nonuniform noise.
@@ II Problem Description
where $\theta_{m}\in(0,\pi)$ represents the DOA of the $m$th source, $\mathbf{a}(\theta_{m})=[1~\cdots~e^{-(N-1)\jmath\pi\cos(\theta_{m})}]^{T}\in\mathbb{C}^{N}$ is the steering vector of the $m$th source, and $\mathbf{A}=[\mathbf{a}(\theta_{1})~\cdots~\mathbf{a}(\theta_{M})]$ is the array manifold matrix. Additionally, $\mathbf{s}(t)=[s_{1}(t)~\cdots~s_{M}(t)]^{T}\in\mathbb{C}^{M}$ is the source signal vector, $\mathbf{v}(t)$ means a spatially nonuniform Gaussian white noise vector with mean zero and covariance $\mathbf{Q}=\mathrm{diag}\{\sigma^{2}_{1},\dots,\sigma^{2}_{N}\}\succ\mathbf{0}_{N}$, i.e., $\mathbf{v}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{Q})$.
@@
where $\boldsymbol{\Phi}=(\boldsymbol{\theta},\mathbf{P},\mathbf{Q})$ with $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{M})$, $p(\cdot)$ is the probability density function, and $\hat{\mathbf{R}}=(1/L)\sum_{t=1}^{L}\mathbf{y}(t)\mathbf{y}^{H}(t)\succ\mathbf{0}_{N}$ [^1] In this letter, we assume $\hat{\mathbf{R}}\succ\mathbf{0}_{N}$, which requires $L\geq N$.. Importantly, we need to solve the equivalent ML direction finding problem [11, 12]:
@@
In order to solve problem (3), the ECME algorithm has been designed in the previous work [13]. In the next section, we will design the ECME algorithm using FA for the subspace based DOA estimation.
III MLFA Based Subspace Estimation
The ML direction finding problem (3) is intractable since $\mathbf{C}=\mathbf{APA}^{H}+\mathbf{Q}$ has excessive constraints on free parameters (unknowns) to be estimated. To alleviate this issue, we should release as many constraints as possible and introduce as many free parameters as possible. Thus, we aim not to solve problem (3) in this letter.
III-A MLFA Estimation Problem
@@
where all elements in $\mathbf{S}\in\mathbb{C}^{N\times M}$ are free parameters (factor loadings) and hence $\mathbf{SS}^{H}$ must be able to equal any $\mathbf{APA}^{H}$. Here, we also use $\mathbf{S}$ to estimate $\mathcal{C}(\mathbf{A})$. Eq. (4) makes problem (3) become the classic MLFA estimation problem [2]:
@@
For solving problem (5), the IMLSE method utilizes a fixed-point iteration and sequentially (not simultaneously) updates the estimates of $\mathbf{S}$ and $\mathbf{Q}$ at each iteration by explicit formulas [7]. However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [8]. Hence, they propose the FAAN method with excellent convergence properties. We outline the FAAN method in the next subsection.
III-B FAAN Method
When a direct minimization over all free parameters is very difficult, an iterative (coordinate descent) approach or alternating minimization can always be adopted [14]. As mentioned before, the FAAN method, composed of two steps at each iteration, has been proposed for problem (5). Specifically, since $\mathbf{C}=\mathbf{SS}^{H}+\mathbf{Q}$ makes updating the estimates of $\mathbf{S}$ and $\mathbf{Q}$ difficult, a new parametrization of $\mathbf{C}$ is designed as
@@ III-B FAAN Method
where $\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}$ and $\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}$ is the eigen-decomposition of $\mathbf{Q}^{-1/2}\mathbf{SS}^{H}\mathbf{Q}^{-1/2}$. Thus, problem (5) is changed to
$\min_{\mathbf{U},\boldsymbol{\Lambda}\succeq\mathbf{0}_{M},\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{Q})+\ln\mathrm{Det}(\boldsymbol{\Lambda}+\mathbf{I}_{M})+$
$\mathrm{Tr}\big{\{}\mathbf{Q}^{-1/2}\hat{\mathbf{R}}\mathbf{Q}^{-1/2}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})^{-1}\big{\}}.$ (7)
According to (III-B), the first step at the $k$th iteration of the FAAN method simultaneously obtains $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ by
$\min_{\mathbf{U},\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}}\ln\mathrm{Det}(\boldsymbol{\Lambda}+\mathbf{I}_{M})+\mathrm{Tr}\big{\{}\tilde{\mathbf{R}}^{(k)}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})^{-1}\big{\}},$ (8)
@@
$\Lambda_{m,m}^{(k)}$ $=$ $(\lambda_{m}^{(k)}-1)_{+},m=1,\dots,M$ (9b)
@@
$\sigma_{n}=\frac{1}{2}\Big{(}b_{n}^{(k)}+\sqrt{[b_{n}^{(k)}]^{2}+4c_{n}^{(k)}}\Big{)}>0,n=1,\dots,N$ (11)
@@
1: Initialize $\mathbf{Q}^{(0)}$, $K_{1}$, $K_{2}$, and $k=1$.
@@
4:  Obtain $\mathbf{Q}^{(k)}$ using (11) $K_{2}$ times.
@@
IV ECME Algorithm
The ECME algorithm has been designed for problem (5) in [6] and yields rapider convergence than the EM algorithm in [15]. However, a CM-step at each iteration utilizes Newton-Raphson to update the estimate of $\mathbf{Q}$ by the actual or negative LLF $f(\mathbf{S},\mathbf{Q})=\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}$ in (5), which leads to that this step is complex and the monotonicity cannot be guaranteed. In this section, we again design the ECME algorithm with excellent convergence.
IV-A Algorithm Procedure
For problem (5), we do not adopt the complete data $(\mathbf{Y},\mathbf{X})$ with $\mathbf{X}=\big{[}\mathbf{x}(1)~\cdots~\mathbf{x}(L)\big{]}$ in [6, 15] and adopt the simpler complete data $(\mathbf{X},\mathbf{V})$ with $\mathbf{V}=\big{[}\mathbf{v}(1)~\cdots~\mathbf{v}(L)\big{]}$ [^3] Numerical results show that the ECME algorithm with $(\mathbf{X},\mathbf{V})$ yields the same convergence as that with $(\mathbf{Y},\mathbf{X})$. However, the ECME algorithm with $(\mathbf{X},\mathbf{V})$ needs fewer computations at each iteration.. Thus, we rewrite the snapshots or measurements $\mathbf{Y}$ in (2) by
@@ IV-A Algorithm Procedure
where $\mathbf{x}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})$ contains $M$ common factors and is uncorrelated with $\mathbf{v}(t)$. Then, the complete-data LLF is
$l(\mathbf{X},\mathbf{V};\mathbf{Q})$ $=$ $\ln p(\mathbf{X})+\ln p(\mathbf{V};\mathbf{Q})$ (13)
@@
where $c$ is a term unrelated to any free parameters in $(\mathbf{S},\mathbf{Q})$, $\hat{\mathbf{R}}_{v}=(1/L)\sum_{t=1}^{L}\mathbf{v}(t)\mathbf{v}^{H}(t)$. With (13), we can construct the ECME algorithm, whose E-step and two CM-steps at the $k$th iteration are derived below.
IV-A1 First CM-step
To avoid providing $\mathbf{S}^{(0)}$, let this step only obtain $\mathbf{S}^{(k)}$ by minimizing $f(\mathbf{S},\mathbf{Q})$ with respect to $\mathbf{S}$, i.e.,
@@ IV-A1 First CM-step
This problem is very classic in MLFA and its classic solution is [2, 3]
$\mathbf{S}^{(k)}=[\mathbf{Q}^{(k-1)}]^{1/2}\mathbf{B}^{(k)}[\boldsymbol{\Lambda}^{(k)}]^{1/2}\in\mathbb{C}^{N\times M},$ (15)
where $\mathbf{B}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ are from (9).
IV-A2 E-step
Given $\mathbf{Y}$, $\mathbf{S}^{(k)}$, and $\mathbf{Q}^{(k-1)}$, this step finds the expectation of (13) by
$l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ $=$ $\mathrm{E}\big{\{}l(\mathbf{X},\mathbf{V};\mathbf{Q})\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\}$ (16)

### v2

@@
The ECME Algorithm Using Factor Analysis for DOA Estimation in Nonuniform Noise
Mingyan Gong
1 Introduction
@@
Factor analysis (FA) has been a tool in psychology for over one hundred years [1] and its purpose is to extract or estimate the low-rank covariance matrix from a measurement covariance matrix. To this end, several techniques have been developed, but the only one available, which provides efficient and consistent estimators, is maximum likelihood FA (MLFA) [2, 3, 4]. Unfortunately, these estimators cannot be expressed in closed form, which leads to a variety of iterative approaches [5]. In particular, one of the most promising is the Expectation-Maximization (EM) algorithm [6].
@@
The EM algorithm regards observations or measurements as the incomplete data and designs the complete data containing missing (latent or unobserved) data. The complete data is so manipulable that its log-likelihood function (LLF) is more tractable than the incomplete-data or actual LLF. Hence, the EM algorithm uses the complete-data LLF to efficiently update the estimates of parameters, which can still increase the actual likelihood. To be specific, each iteration of the EM algorithm consists of an expectation step (E-step) and a maximization step (M-step). However, the M-step requires maximizing the expected complete-data LLF provided by the E-step, which causes two shortcomings: 1) slow convergence, and 2) difficult maximization in numerous complex cases. To make up for both shortcomings, the Expectation/Conditional Maximization Either (ECME) algorithm, an extension of the EM algorithm, has been proposed [7] and later applied to MLFA [8]. However, a conditional M-step (CM-step) at its each iteration utilizes Newton-Raphson, instead of an explicit formula, to update the estimates of partial parameters by the actual LLF and thus is complex as well as unstable.
@@
In addition to psychometrics, econometrics, statistics and so on, FA recently finds applications in array signal processing, especially including direction of arrival (DOA) estimation in unknown nonuniform noise [9, 10]. The nonuniform noise assumption can let the signal received at an array be modeled by FA so that MLFA is used to estimate the signal subspace for the subspace based DOA estimation [11]. For this purpose, an iterative approach, called Iterative ML Subspace Estimation (IMLSE), is proposed in [9]. This method sequentially updates the estimates of parameters at each iteration by explicit formulas and is computationally attractive. However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [10]. Afterwards, they propose a novel iterative approach called Factor Analysis for Anisotropic Noise (FAAN), which has excellent convergence properties.
@@
In this article, the ECME algorithm is designed again for MLFA, which has almost the same computational complexity at each iteration as the FAAN method. However, numerical results show that the ECME algorithm yields faster stable convergence and is computationally more efficient. Importantly, the signal subspace estimated by the ECME algorithm can be employed for the subspace based DOA estimation in unknown nonuniform noise.
@@ 2 Problem Description
where $\theta_{m}\in(0,\pi)$ represents the DOA of the $m$th source signal, $\mathbf{a}(\theta_{m})=[1~\cdots~e^{-(N-1)\jmath\pi\cos(\theta_{m})}]^{T}\in\mathbb{C}^{N}$ is the steering vector of the $m$th source, and $\mathbf{A}=[\mathbf{a}(\theta_{1})~\cdots~\mathbf{a}(\theta_{M})]$ is the array manifold matrix. Additionally, $\mathbf{s}(t)=[s_{1}(t)~\cdots~s_{M}(t)]^{T}\in\mathbb{C}^{M}$ is the source signal vector, $\mathbf{v}(t)$ is a spatially nonuniform Gaussian white noise vector with mean zero and diagonal covariance matrix $\mathbf{Q}=\mathrm{diag}\{\sigma^{2}_{1},\dots,\sigma^{2}_{N}\}\succ\mathbf{0}_{N}$, i.e., $\mathbf{v}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{Q})$.
@@
where $\boldsymbol{\Phi}=(\boldsymbol{\theta},\mathbf{P},\mathbf{Q})$ with $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{M})$, $p(\cdot)$ is the probability density function, and $\hat{\mathbf{R}}=(1/L)\sum_{t=1}^{L}\mathbf{y}(t)\mathbf{y}^{H}(t)\succ\mathbf{0}_{N}$ [^1] In this article, we assume $\hat{\mathbf{R}}\succ\mathbf{0}_{N}$, which requires $L\geq N$.. Importantly, we need to solve the equivalent ML direction finding problem [13, 14]:
@@
In order to solve problem (3), the ECME algorithm has been designed in the previous work [15]. In next sections, we will design the ECME algorithm using FA for the subspace based DOA estimation.
3 Subspace Estimation Using MLFA
The ML direction finding problem (3) is intractable since $\mathbf{C}=\mathbf{APA}^{H}+\mathbf{Q}$ has excessive constraints on free parameters (unknowns) to be estimated. To alleviate this issue, we should release as many constraints as possible and introduce as many free parameters as possible. Thus, we aim not to solve problem (3) in this article.
3.1 MLFA Estimation Problem
@@
where all elements in $\mathbf{S}\in\mathbb{C}^{N\times M}$ are free parameters (factor loadings) and hence $\mathbf{SS}^{H}$ must be able to equal any $\mathbf{APA}^{H}$. Here, we also use $\mathbf{S}$ to estimate $\mathcal{C}(\mathbf{A})$. Eq. (4) makes problem (3) become the classic MLFA estimation problem [2, 4, 3, 5]:
@@
For solving problem (5), the IMLSE method utilizes a fixed-point iteration and sequentially (not simultaneously) updates the estimates of $\mathbf{S}$ and $\mathbf{Q}$ at each iteration by two explicit formulas [9]. However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [10]. Hence, they propose the FAAN method with excellent convergence properties. We outline the FAAN method in the next subsection.
3.2 FAAN Method
When a direct minimization over all free parameters is very difficult, an iterative (coordinate descent) approach or alternating minimization can always be adopted [16]. As mentioned before, the FAAN method has been proposed for problem (5) and its each iteration consists of two steps. Specifically, since $\mathbf{C}=\mathbf{SS}^{H}+\mathbf{Q}$ makes updating the estimates of $\mathbf{S}$ and $\mathbf{Q}$ difficult, a new parametrization of $\mathbf{C}$ is designed as
@@ 3.2 FAAN Method
where $\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}$, $\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}$ is the eigen-decomposition of $\mathbf{Q}^{-1/2}\mathbf{SS}^{H}\mathbf{Q}^{-1/2}$, and $\mathbf{S}=\mathbf{Q}^{1/2}\mathbf{U}\boldsymbol{\Lambda}^{1/2}$. Thus, problem (5) is changed to
$\min_{\mathbf{U}^{H}\mathbf{U}=\mathbf{I}_{M},\boldsymbol{\Lambda}\succeq\mathbf{0}_{M},\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{Q})+\ln\mathrm{Det}(\boldsymbol{\Lambda}+\mathbf{I}_{M})$
$+\mathrm{Tr}\big{\{}\mathbf{Q}^{-1/2}\hat{\mathbf{R}}\mathbf{Q}^{-1/2}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})^{-1}\big{\}}.$ (7)
According to (3.2), the first step at the $k$th iteration of the FAAN method simultaneously obtains $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ by
$\min_{\mathbf{U}^{H}\mathbf{U}=\mathbf{I}_{M},\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}}\ln\mathrm{Det}(\boldsymbol{\Lambda}+\mathbf{I}_{M})+$
$\mathrm{Tr}\big{\{}\tilde{\mathbf{R}}^{(k)}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})^{-1}\big{\}},$ (8)
@@
$\Lambda_{m,m}^{(k)}$ $=$ $(\lambda_{m}^{(k)}-1)_{+}\geq 0,m=1,\dots,M$ (9b)
@@
$\sigma_{n}$ $=$ $\frac{1}{2}\Big{(}b_{n}^{(k)}+\sqrt{[b_{n}^{(k)}]^{2}+4c_{n}^{(k)}}\Big{)}>0,n=1,\dots,N$
@@
1: Initialize $\mathbf{Q}^{(0)}\succ\mathbf{0}_{N}$, $K_{1}$, $K_{2}$, and $k=1$.
@@
4:  Obtain $\mathbf{Q}^{(k)}$ using (3.2) $K_{2}$ times.
@@
4 ECME Algorithm
The ECME algorithm has been designed for problem (5) in [8] and yields rapider convergence than the EM algorithm in [17]. However, a CM-step at its each iteration utilizes Newton-Raphson, instead of an explicit formula, to update the estimate of $\mathbf{Q}$ by the actual or negative LLF $f(\mathbf{S},\mathbf{Q})=\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}$ in (5), leading to that this step is complex and the monotonicity cannot be guaranteed. In this section, we again design the ECME algorithm with excellent convergence.
4.1 Algorithm Procedure
For problem (5), we do not adopt the complete data $(\mathbf{Y},\mathbf{X})$ with $\mathbf{X}=\big{[}\mathbf{x}(1)~\cdots~\mathbf{x}(L)\big{]}$ in [8, 17] and adopt the simpler complete data $(\mathbf{X},\mathbf{V})$ with $\mathbf{V}=\big{[}\mathbf{v}(1)~\cdots~\mathbf{v}(L)\big{]}$ [^3] Numerical results show that the ECME algorithm with $(\mathbf{X},\mathbf{V})$ yields the same convergence as that with $(\mathbf{Y},\mathbf{X})$. However, the ECME algorithm with $(\mathbf{X},\mathbf{V})$ needs fewer computations at each iteration.. Thus, we rewrite the snapshots or measurements $\mathbf{Y}$ in (2) by [2, 4, 3]
@@ 4.1 Algorithm Procedure
where $\mathbf{x}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})$ is composed of $M$ common factors and uncorrelated with $\mathbf{v}(t)$. Then, the complete-data LLF is
$l(\mathbf{X},\mathbf{V};\mathbf{Q})$ $=$ $\ln p(\mathbf{X})+\ln p(\mathbf{V};\mathbf{Q})$
@@
where $c$ is a term unrelated to any free parameters in $(\mathbf{S},\mathbf{Q})$, $\hat{\mathbf{R}}_{v}=(1/L)\sum_{t=1}^{L}\mathbf{v}(t)\mathbf{v}^{H}(t)$. With (4.1), we can construct the ECME algorithm, whose E-step and two CM-steps at the $k$th iteration are derived below.
4.1.1 First CM-step:
Given $\mathbf{Q}=\mathbf{Q}^{(k-1)}$, let this step only obtain $\mathbf{S}^{(k)}$ by minimizing the actual or negative LLF $f(\mathbf{S},\mathbf{Q})$ with respect to $\mathbf{S}$, i.e.,
@@ 4.1.1 First CM-step:
This problem is classic [2, 4, 3] but the classic solutions in [18, 19] may not exist. Thus, we resort to the FAAN method by
$\mathbf{S}^{(k)}=[\mathbf{Q}^{(k-1)}]^{1/2}\mathbf{U}^{(k)}[\boldsymbol{\Lambda}^{(k)}]^{1/2}\in\mathbb{C}^{N\times M},$ (15)
where $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ are from (9).
4.1.2 E-step:
Given $\mathbf{Y}$, $\mathbf{S}^{(k)}$, and $\mathbf{Q}^{(k-1)}$, this step finds the expectation of $l(\mathbf{X},\mathbf{V};\mathbf{Q})$ in (4.1) by
$l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ $=$ $\mathrm{E}\big{\{}l(\mathbf{X},\mathbf{V};\mathbf{Q})\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\}$